The following is the balance sheet of Korver Supply Company at December 31, 2017.

KORVER SUPPLY COMPANY
Balance Sheet
At December 31, 2017
Assets
Cash $ 160,000
Accounts receivable 320,000
Inventories 270,000
Furniture and fixtures, net 180,000
Total assets $ 930,000
Liabilities and Shareholders’ Equity
Accounts payable (for merchandise) $ 270,000
Note payable 280,000
Interest payable 7,000
Common stock 110,000
Retained earnings 263,000
Total liabilities and shareholders' equity $ 930,000


Transactions during 2018 were as follows:

1. Sales to customers on account $ 930,000
2. Cash collected from customers 910,000
3. Purchase of merchandise on account 620,000
4. Cash payment to suppliers 630,000
5. Cost of merchandise sold 570,000
6. Cash paid for operating expenses 290,000
7. Cash paid for interest on note 14,000


The note payable is dated June 30, 2017 and is due on June 30, 2019. Interest at 5% is payable annually on June 30. Depreciation on the furniture and fixtures for the year is $33,000. The furniture and fixtures originally cost $430,000.

Required:
Prepare a classified balance sheet at December 31, 2018 (ignore income taxes). (Amounts to be deducted should be indicated by a minus sign.)
Answer #1

Income statement (2018)

Sales revenue 930,000
Cost of merchandise sold - 570,000
Gross profit 360,000
Cash Operating expenses - 290,000
Depreciation expense - 33,000
Operating profit 37,000
Interest expense - 14,000
Net income 23,000

Retained earnings, ending = Retained earnings, beginning + Net income

= 263,000 + 23,000

= $286,000

Ending cash balance = Beginning cash balance + Cash collected from customers - Cash payment to suppliers - Cash paid for operating expenses - Cash paid for interest on note

= 160,000 + 910,000 - 630,000 - 290,000 - 14,000

= $136,000

Accounts receivable ending balance = Accounts receivable beginning + Sales to customers on account- Cash collected from customers

= 320,000 + 930,000 - 910,000

= $340,000

Inventories, ending balance = Inventories, beginning + Purchase of merchandise on account - Cost of merchandise sold

= 270,000 + 620,000 - 570,000

= $320,000

Accounts payable, ending balance = Accounts payable, beginning + Purchase of merchandise on account - Cash payment to suppliers

= 270,000 + 620,000 - 630,000

= $260,000

Furniture and fixtures, ending balance = Furniture and fixtures, beginning balance - Depreciation expense

= 180,000 - 33,000

= $147,000

Balance sheet of Korver Supply Company at December 31, 2018.

KORVER SUPPLY COMPANY
Balance Sheet
At December 31, 2018
Assets
Cash $ 136,000
Accounts receivable 340,000
Inventories 320,000
Furniture and fixtures, net 147,000
Total assets $ 943,000
Liabilities and Shareholders’ Equity
Accounts payable (for merchandise) $ 260,000
Note payable 280,000
Interest payable 7,000
Common stock 110,000
Retained earnings 286,000
Total liabilities and shareholders' equity $ 943,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.